### v4

# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

Abstract.
Generative retrieval methods utilize generative sequential modeling techniques, such as transformers, to generate candidate items for recommender systems. These methods have demonstrated promising results in academic benchmarks, surpassing traditional retrieval models like two-tower architectures. However, current generative retrieval methods lack the scalability required for industrial recommender systems, and they are insufficiently flexible to satisfy the multiple metric requirements of modern systems.
This paper introduces PinRec, a novel generative retrieval model developed for applications at Pinterest. PinRec utilizes outcome-conditioned generation, enabling modelers to specify how to balance various outcome metrics, such as the number of saves and clicks, to effectively align with business goals and user exploration. Additionally, PinRec incorporates multi-token generation to enhance output diversity while optimizing generation. Our experiments demonstrate that PinRec can successfully balance performance, diversity, and efficiency, delivering a significant positive impact such as +2% sitewide clicks and +4% search repins. This paper marks a significant milestone in generative retrieval, as it presents, to our knowledge, the first rigorous study on implementing generative retrieval at the scale of Pinterest.
generative retrieval, recommender systems, transformer, item recommendation
journalyear: 2025copyright: ccccs: Information systems Retrieval models and rankingccs: Information systems Personalization
Figure 1. Illustration of PinRec, a generative item retrieval technique for heterogeneous user journeys on Pinterest. Sequences of user searches, engagements, and outcome-conditioning (bottom) are used to recommend Pins (top).
1. Introduction
Pinterest is a visual discovery platform where over 550 million monthly active users search for, save, and shop for ideas, represented by Pins (pins). Given the massive volume of Pins — for instance, users save over 1.5 billion Pins per week (pins2) — it is essential to provide personalized recommendations in Pinterest feeds and search results. To achieve this, recommender systems often leverage a two-stage approach: retrieval and ranking. Retrieval models select thousands of potentially relevant candidate items, then ranking models prioritize the top items to display to the user. Since ranking models depend on the retrieval results, it is important for retrieval models to identify diverse, fresh, and relevant items.
Traditionally, retrieval methods employ “two-tower” models. The query tower computes query embeddings, while the item tower computes item embeddings. These models retrieve items with the highest similarity scores given a query embedding. As an alternative to two-tower models, recent work has proposed generative retrieval. Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (vaswani2017attention), and then use those tokens to retrieve items. Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (KangM18; Recformer23; UnifyingGenRetrieval24).
Using generative retrieval models for large scale recommender systems poses many challenges at the scale of Pinterest. Firstly, the generative model should be capable of efficiently optimizing for different engagement outcomes such as maximizing social actions like saves and comments versus passive engagements like clicks, because business needs are constantly evolving to optimize combinations of different metrics. Secondly, the generated item sequences must possess sufficient diversity to facilitate users’ exploration of varied ideas; one-step generation of a single item representation may not suffice. Thirdly, given their increased complexity and sequential generation, generating a set of recommended items requires higher computational cost than computing an embedding with two-tower models.
To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges. Most methods developed within academic settings (KangM18; UniS4Rec22; Recformer23) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic. While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (HSTU24; Tiger23) and are not capable of optimizing multiple target outcomes. Moreover, these methods do not address the challenge of generating a sequence of diverse items. Because generating multiple items sequentially is computationally expensive, existing approaches typically generate only the next item rather than a full set at once.
Figure 2. Modeling diagram of PinRec
In this paper, we present PinRec, a generative retrieval technique for real-world recommender systems (Figure 2). PinRec builds on other generative retrieval models, leveraging transformers to perform next token prediction (KangM18; HSTU24; Tiger23). Unlike previous methods, however, PinRec addresses and resolves three aforementioned challenges for generation of optimal recommendations: multi-objective generative capabilities, computational efficiency, and item diversity. Our main contributions in the paper are summarized below.
  • Outcome-Conditioned Generative Retrieval: PinRec employs outcome-conditioned generation, enabling the optimization of multiple business metrics. For each outcome (e.g., business metrics such as user saves or clicks), PinRec generates the next item token that is most likely to maximize that specific outcome, and then retrieves candidate items to balance these objectives according to the desired trade-offs (Fig. 2). Through both offline and online evaluations, we show that tailoring item recommendations to particular outcomes enables designers to effectively achieve those specified objectives.
  • Multi-Token Generation for Efficient Sequence Generation: One obstacle to generating an entire sequence of items for retrieval is that traditional models typically produce only one item at a time. To address this, we propose a multi-token generation framework (Multitoken24), in which PinRec generates multiple future items in parallel, as illustrated in Figure 2. This approach significantly reduces the number of steps needed to produce diverse output embeddings, resulting in substantial offline and online improvements in performance, diversity, and efficiency.
  • Efficient Serving: We detail how we optimize autoregressive generation for optimal item recommendation, implement the serving infrastructure, and allow for horizontal scaling to serve hundreds of millions of users. To our knowledge, this study is the first to disclose the details of infrastructure optimization for industry-scale generative retrieval systems.
We will make the model code publicly available upon publication.
2. Related Work
Sequential Recommendation Systems: Sequential recommendation systems operate on a sequences of items with which a user has interacted (KangM18). Earlier work in academia has used recurrent neural networks (RNN) (GRU4REC16), but later work has often adopted transformers (KangM18; BERT4REC19). Recent studies have harnessed the extensive knowledge of pre-trained language models (UniS4Rec22; Recformer23).
In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pinnerformer22). The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models) (Spotify20; rethink; Transact23; TWINV224). Another approach in sequential recommendation systems is the generative approach, where the model learns to generate the items to recommend. Tiger23 used transformers to generate items by producing a sequence of semantic ID tokens (RQVAE22; SemanticIDGeneral24). HSTU24 also proposed a generative retrieval method. However, these studies have mainly focused on generating a single item to optimize a single action. In contrast, our research aims to generate sequences of multiple items to optimize for multiple outcomes.
Multi-Objective Generation: We discuss generative models that consider multiple objectives. While RL algorithms have shown this capability in industry (SelfsupervisedRL20), studies in language modeling suggest RL often needs extensive tuning for optimal generation (PPO22). Recent research has introduced more streamlined and scalable techniques based on supervised learning, such as goal-conditioned methods (IterativeSupervised21) and prompting strategies (PromptingRL22). PinRec builds on these approaches to enhance scalability.
3. Problem Setup
We formulate the generative sequential recommendation task as follows. Consider a system with users $\mathcal{U}$, items $\mathcal{I}$, actions $\mathcal{A}$, and surfaces $\mathcal{S}$. Our objective is to generate an ordered list of $N$ item recommendations $R(u,t)\subset\mathcal{I}$ for user $u\in\mathcal{U}$ at timestep $t$.
Item Types and Actions. Each item $i\in\mathcal{I}$ belongs to an item type $T(i)\in\mathcal{T}$, where $\mathcal{T}=\{T(i)\mid i\in\mathcal{I}\}$ represents the set of all item types in the system (e.g., Pins, boards, search queries). Users can perform various actions $\mathcal{A}=\{\text{click},\text{save},\text{share},\text{comment},\ldots\}$ on these items across different surfaces. Each interaction occurs on a specific surface $s\in\mathcal{S}$, where surfaces represent different recommendation products, such as homefeed, search results, board pages, and related pin feed.
User History. For each user $u$, we maintain their interaction history $H(u,t_{\text{max}})=\langle(i_{1},a_{1},s_{1},t_{1}),(i_{2},a_{2},s_{2},t_{2}),\ldots,(i_{m},a_{m},s_{m},t_{m})\rangle$, where each tuple $(i_{j},a_{j},s_{j},t_{j})$ represents an interaction with item $i_{j}\in\mathcal{I}$ via action $a_{j}\in\mathcal{A}$ on surface $s_{j}\in\mathcal{S}$ at timestep $t_{j}\leq t_{\text{max}}$. The sequence is chronologically ordered such that $t_{1}\leq t_{2}\leq\ldots\leq t_{m}$.
Sequential Recommendation Model. We propose a parameterized sequence model $f_{\theta}:\mathcal{H}\times\mathcal{T}\times\mathcal{S}\rightarrow\mathbb{R}^{|\mathcal{I}|}$ that maps a user’s interaction history, item types, and surface context to item scores, where $\mathcal{H}$ denotes the space of all possible interaction histories. The model parameters $\theta$ are optimized over the collection of user interaction sequences $\{H(u,t_{\text{max}})\mid u\in\mathcal{U}\}$ to maximize the likelihood of observed user interactions while accounting for the heterogeneous nature of actions and surfaces. This enables the model to learn nuanced behavioral patterns that depend on both what users interact with and how and where they interact with it.
4. PinRec
In this section, we describe the model architecture and serving infrastructure for PinRec. We start with an overview of the model, including outcome-conditioned generation and multi-token prediction. Next, we explain how items in the sequence are encoded and processed, and how the training loss is computed. Finally, we describe inference mode operation and provide details about our real-time serving architecture, which is powered by Nvidia Triton.
4.1. Model Architecture
We use a transformer decoder architecture as the backbone of our PinRec model. The user interaction history is first featurized and fed into a stack of transformer layers along with learned position embeddings to capture temporal ordering. At each layer, we apply causal attention masking to ensure that tokens at time $t$ can only attend to tokens at positions $t^{\prime}\leq t$, preventing information leakage from future interactions. The stack of $L$ transformer layers outputs a sequence of contextualized hidden states: $\mathbf{h}_{u,t}=\text{TransformerStack}^{(L)}(\mathbf{x}_{u,1:t})$.
where $\mathbf{x}_{u,1:t}$ represents the featurized user interaction sequence up to timestep $t$, and $\mathbf{h}_{u,t}$ is the resulting hidden state at position $t$ from the final layer. These hidden states are then processed through an output head $O$ to generate the next item predictions.
Outcome-Conditioned Generation. Unlike language modeling, where the output sequence receives a single reward value (PPO22), users interact with recommended items through diverse engagement types—saving, sharing, commenting, or clicking. To effectively optimize across these varied engagement metrics, generative retrieval methods must account for this behavioral diversity (MultiBehaviorSASRec24). Furthermore, practical deployment often requires specifying the target surface or context for the next recommendation, even when the user’s historical sequence spans multiple surfaces.
To address these requirements, we design our model to accept conditioning signals at each generation step, enabling dynamic control over model behavior during inference. Specifically, we condition the output head $O$ on a desired set of outcomes, allowing the model to generate targeted recommendations. We implement this through learnable embeddings for each possible outcome type (e.g., engagement actions, surface constraints). By conditioning the generation process on these outcome representations, our model produces outcome-specific, personalized item representations:
(1) $\hat{\mathbf{i}}_{u,t}=O(\mathbf{h}_{u,t},\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k})$
where $\mathbf{h}_{u,t}$ is the hidden state at timestep $t$, $\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k}$ are the conditioning embeddings for different outcome types (e.g., $\mathbf{c}_{1}=\mathbf{e}_{\text{action}}$, $\mathbf{c}_{2}=\mathbf{e}_{\text{surface}}$), and $\hat{\mathbf{i}}_{u,t}$ is the resulting conditioned item representation.
Temporal Multi-token Prediction. Multi-token prediction in language models involves predicting multiple successive tokens at positions $t+1$ to $t+k$ given the current position $t$ (gloeckle2024better). This is typically implemented using separate output heads for each future position. However, in recommender systems, previous research (Pinnerformer22) suggests that the difference in time between actions is much more important than the position index alone, because a single position increment can correspond to very different time intervals for active users versus dormant users. To address this, we propose a time-based multi-token approach, in which tokens are generated for each time interval $\delta$ (e.g., generating recommendations for the user for the present, 30 seconds later, 1 minute later, and so on). Specifically, we condition our output head on the relative temporal distance from the current action, enabling a single parameterized head to generate predictions at various future time intervals:
(2) $\hat{\mathbf{i}}_{u,t+\delta}=O(\mathbf{h}_{u,t},\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k},\mathbf{e}_{\delta})$
where $\mathbf{e}_{\delta}$ is the learnable embedding for temporal offset $\delta\in\{1,2,\ldots,k\}$, and $\hat{\mathbf{i}}_{u,t+\delta}$ represents the predicted item at timestep $t+\delta$.
4.2. Input Sequence Representation
Given the heterogeneous nature of user interactions $(i_{j},a_{j},s_{j},t_{j})$ in our interaction history $H(u,t_{\text{max}})$, we need to transform each element into appropriate representations for the transformer architecture. A crucial design decision is defining representations for heterogeneous item types. There are two main ways to represent items in a sequence: (1) by using dense embedding vectors (HSTU24; KangM18), and (2) by learning a sequence of discrete IDs for each item, known as Semantic IDs (Tiger23; SemanticIDGeneral24). We explored both approaches and ultimately chose to use dense embedding vectors. In our experiments, we observed that Semantic IDs frequently suffered from representational collapse—many items were assigned the same Semantic ID, as also noted in (MultiBehaviorSASRec24; Hourglass24)—and consistently produced lower evaluation metrics (see Table 1(UnifyingGenRetrieval24). While optimizing performance with Semantic IDs remains an interesting direction for future work, our core contributions are representation-agnostic.
We now detail how each component of the interaction tuple $(i_{j},a_{j},s_{j},t_{j})$ is processed and integrated into our model architecture.
Item Embeddings. We generate dense embeddings for different entity types ($\tau\in\mathcal{T}$) using type-specific embedders that are learned end-to-end with the model. Each item type $\tau$ has a dedicated embedder $f_{\tau}$ that maps type-specific features to the shared representation space. While we describe embedders for Pins and search queries in this work, the approach can be extended to other entity types such as boards, users, or topics.
Pin Embeddings. For Pins, we combine semantic embedding features with learned Id embeddings (10.1145/3640457.3688053). The semantic representation uses the pre-trained OmniSage embedding (badrinath2025omnisage), which captures visual and textual content along with the engagement via the Pin-board graph. Additionally, we implement a large-scale ID embedding system where each Pin identifier is hashed into $k$ different hash functions, each mapping to embedding tables of size $d/k$. These $k$ embeddings are concatenated to form a $d$-dimensional ID embedding, providing a distributed representation that scales efficiently. The ID embeddings constitute sparse parameters that allow us to scale model capacity with minimal impact on serving latency and computational cost, as only a small subset of embeddings are accessed for any given batch of items. The concatenated OmniSage and ID embeddings are passed through 2-3 fully connected layers with ReLU activation and layer normalization, followed by $L_{2}$ normalization to produce the final Pin embedding $\mathbf{i}_{\text{pin}}$.
Query Embeddings. For search queries, we use only semantic representations as ID embeddings have not provided value in offline metrics for this entity type. We leverage the pre-trained OmniSearchSage query representation (agarwal2024omnisearchsage) which captures semantic query understanding. This representation is passed through a similar MLP architecture (2-3 fully connected layers with ReLU activation and layer normalization) followed by $L_{2}$ normalization to generate the final query embedding $\mathbf{i}_{\text{query}}$.
Temporal Embeddings. For each interaction timestep $t_{j}$, we extract absolute timestamp and time since previous action, capturing both global temporal context and local session dynamics. Our encoder applies sinusoidal transformations at multiple frequencies. For absolute timestamps, we use predefined periods $P=\{p_{1},p_{2},\ldots,p_{k}\}$ to capture daily, weekly, and seasonal patterns: $\sin(t_{\text{abs}}/p_{i}),\cos(t_{\text{abs}}/p_{i})$. For relative time differences, we use log-scale frequencies augmented with learnable phase parameters.
Sequence Construction. The final input sequence to the transformer is constructed by concatenating item and temporal embeddings, then adding surface embeddings for each interaction. This creates a rich contextual representation that captures what users interacted with (item), when they interacted (temporal), and where the interaction occurred (surface) within the platform.
4.3. Training Objective
During training, we employ teacher forcing with a next-item prediction objective. The hidden states $\mathbf{h}_{u,t}$ are processed through the conditioned output head $O$ to generate item representations that predict future items in the sequence. For outcome-conditioned generation, we condition the output head on the actual values of the target item during training. For multi-token prediction, we sample $k$ positions from the next $N$ positions in the sequence to serve as prediction targets.
Following Pinnerformer22, we compute similarity between the predicted item representation $\hat{\mathbf{i}}_{u,t}$ and candidate item $\mathbf{i}_{c}$ as $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{i}_{c})$, where $Q(\mathbf{i}_{c})$ is a bias correction term estimated using count-min sketches (cormode2005improved) to account for the non-uniform sampling of in-batch negatives.
The sampled softmax loss maximizes similarity to target items relative to negative samples $\mathbf{N}$:
(3) $L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}})=-\log\frac{\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}}))}{\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}}))+\sum_{\mathbf{i}_{n}\in\mathbf{N}}\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{n}))}$
We average this loss over all positions in each user sequence. For standard next-token prediction, the target is simply the immediate next item:
(4) $L_{\theta}^{\text{next}}(H(u,t_{\max}))=\frac{1}{|H(u,t_{\max})|}\sum_{t=1}^{t_{\max}-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
However, this strict sequential assumption may not align well with user interaction patterns. User interaction sequences differ fundamentally from language: while they exhibit global structure reflecting long-term interests, they lack the local structural constraints found in natural language. On social media platforms, users engage with feed suggestions in arbitrary order within sessions, making strict positional dependencies less meaningful than in natural language modeling.
To better accommodate these interaction patterns, we introduce an intra-feed relaxation that considers all items engaged within the same feed session as valid targets. Given a feed session $F_{s}=\{i_{t_{1}},i_{t_{2}},\ldots,i_{t_{n}}\}$ where $t_{1}<t_{2}<\ldots<t_{n}$, we define the relaxed target set for position $t_{j}$ as:
$T_{t_{j}}^{\text{feed}}=\{i_{t_{k}}\mid t_{k}\in F_{s},t_{k}>t_{j}\}$
This enables a relaxed objective that optimizes for predicting any future item within the same session:
(5) $L_{\theta}^{\text{feed}}(H(u,t_{\mathrm{max}}))=\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t_{j}}\min_{i\in T_{t_{j}}^{\text{feed}}}L_{s}(\hat{\mathbf{i}}_{u,t_{j}},i)$
where we minimize the loss to the best-matching item among all future items within the same feed session, rather than strictly requiring prediction of the immediate next item. Intuitively, this optimization can be posed as: can we predict any of the future items that the user will engage with within the same feed session? We find that this approach is more naturally aligned with user interaction patterns on Pinterest compared to strict next-token prediction.
Finally, we minimize the expected loss across all user sequences in dataset $\mathcal{D}$:
(6) $\arg\min_{\theta}\mathbb{E}_{H(u,t_{\max})\sim\mathcal{D}}[L_{\theta}^{\text{feed}}(H(u,t_{\max}))].$
4.4. Inference
Our PinRec model serves as a candidate generator within the broader recommendation stack, producing diverse item embeddings. The model operates in an autoregressive fashion to generate multiple candidate embeddings, which are then used to retrieve actual items through approximate nearest neighbor search.
To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer. The specific generation process varies based on the model configuration:
Unconditional Generation. For the baseline model, generation is straightforward: at each step, the model produces a single item embedding, which is appended to the sequence along with the target surface embedding and a predefined temporal offset (e.g., 10 seconds). This updated sequence is then used to generate the next embedding.
Outcome-Conditioned Generation. For models with outcome conditioning, for each outcome we generate a candidate embedding, producing multiple candidate embeddings per generation step. For autoregressive continuation, we sample a value for each condition according to the requested distribution over each condition $D_{1},D_{2},\ldots,D_{n}$ and use the corresponding embedding as input for the next step (Fig. 2).
Temporal Multi-Token Generation. When using temporal multi-token prediction, we sample time deltas up to a specified maximum and generate multiple embeddings at different temporal offsets, combined with any specified outcome conditions. For the next autoregressive input, we use the embedding corresponding to the immediate next token.
Across $n$ rollouts, this process generates varying numbers of embeddings depending on the model configuration and input settings. The generated embeddings are then used for candidate retrieval through Approximate Nearest Neighbor (ANN) search.
Budget Allocation. To retrieve $N$ total items, we first allocate retrieval budget across all generated embeddings. For unconditional models, the budget is distributed equally among all embeddings. For outcome-conditioned models, allocation is performed independently for each action type according to the specified budget constraints $\mathcal{B}$, ensuring that each condition receives its designated proportion of the total retrieval budget.
Embedding Compression. Given that autoregressive generation produces multiple output representations, we observe that many embeddings are often similar to one another. During ANN retrieval, these similar embeddings would collapse into overlapping sets of retrieved items, reducing diversity. To address this, we implement an embedding compression strategy: each generated embedding is merged into a previously uncompressed embedding (in generation order) if their cosine similarity exceeds a predefined threshold. When embeddings are compressed together, their allocated budgets are summed to determine the total retrieval budget for the compressed group.
This candidate generation process enables PinRec to serve as a flexible first-stage retrieval system within Pinterest’s recommendation infrastructure. The generated candidates can be further refined by downstream ranking models, allowing the system to benefit from both the generative model’s ability to capture complex user patterns and the precision of specialized ranking algorithms. The outcome-conditioning and multi-token capabilities make the system particularly suitable for multi-objective optimization scenarios common in production recommendation systems.
4.5. Serving
Figure 3. Serving flow for the PinRec system. Green boxes represent services, blue rectangles represent transmitted data, purple cylinders represent indexed data, and beige boxes are steps within the PinRec NVIDIA Triton ensemble.
PinRec is implemented with an in-house, optimized implementation of transformer and served with NVIDIA Triton Inference Server’s ensemble setting and Python backend to perform autoregressive generative retrieval. This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with learned ID embeddings, autoregressive inference for item generation, and Faiss retrieval (Figure 3).
4.5.1. Signal Fetching
Our Signal Service uses a Lambda architecture that combines both batch and real-time processing. The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, storing the sequences and features for each user ID. The real-time component is backed by a RocksDB key-value store (kvstore) that efficiently retrieves user engagements occurring after the batch signal’s cutoff time. The batch and real-time signals are deduplicated and returned to PinRec. On surfaces with context (e.g., Search, Related Pins), model inference also includes the context (e.g., search query, closeup Pin) issued by the user at request time as the final element in the sequence. To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision and later dequantized to FP16 for inference.
4.5.2. Learned ID Embeddings
As discussed in Section 4.2, our pin embedder includes an ID embedding table that maps each pin ID’s hash to an embedding, followed by MLP layers. Since this ID embedder requires significant memory, especially when replicated, it runs as a separate CPU memory-optimized service to avoid competing with the main model for GPU resources. It is served with torchscript to achieve single-digit millisecond latencies.
4.5.3. Autoregressive Inference
We serve the main PinRec model for autoregressive generation on NVIDIA L40S GPUs. To generate an ordered set of item recommendations efficiently, we employ several optimizations for latency and cost efficiency:
CUDA Graphs.
To limit overhead from kernel launches and include other in-built PyTorch optimizations, we use CUDA Graphs with torch.compile to maximally compile the main components of the model.
KV Cache.
During the prefill stage of the transformer, the $K$ and $V$ tensors for the user’s known sequence are cached in a per-batch-request KV Cache. On subsequent decode steps, these are reused, and only new sequence elements are processed to produce $K$ and $V$ for the additional sequence dimension.
4.5.4. Faiss Retrieval
(a) Recent user history specified to produce results in Subfigure 4(b).
(b) Results for the user sequence specified in Subfigure 4(a).
Figure 4. Item recommendations from PinRec-OC for a user history with cat-related search queries and pin interactions (history shown in reverse chronological order). Additional examples are provided in Section A.2.
Our nearest-neighbor retrieval involves a Faiss IVF-HNSW index running on CPU hosts, with inner product as the similarity metric. As with the main model, this is also served with NVIDIA Triton with batching enabled. We run kkNN retrieval to fetch the allocated number of Pins for each embedding, with an overfetch factor for later de-duplication.
Sample results from this complete serving pipeline are visualized for a manually specified user history in Figure 4.
5. Offline Evaluation
In this section, we dissect the performance of the proposed generative recommendation system, PinRec, on generative retrieval across a range of offline evaluations. For the results presented, PinRec uses a 12-layer transformer model (12 attention heads, 768 hidden dimensions) with an 80M-row ID embedding table (256 total dimensions, 8 dimensions per hash). Our experimental setting includes interactions on three major surfaces for the Pinterest platform, namely Homefeed, Related Pins, and Search. We demonstrate that outcome-conditioned generation allows for improved controllability and that multi-token generation leads to significantly improved diversity, performance, and efficiency.
5.1. Offline Evaluation Metrics
We construct the evaluation set using interactions from users disjoint from the training set and evaluate whether the proposed technique can produce representations that retrieve future item embeddings from a set of $1M$ random items (the ”index”). For such retrieval tasks, recall@$k$ is a standard performance metric, measuring the proportion of target items correctly retrieved within the top-$k$ ranked items.
We evaluate model performance using a sequence generalization of recall@$k$ that accounts for the multi-embedding nature of our autoregressive generation. Unlike standard recall, which compares individual predictions to individual targets, our unordered recall metric compares sets of predictions to sets of targets for each user.
Given output item representations $\hat{\mathbf{I}}_{u}=\{\hat{\mathbf{i}}_{u,1},\hat{\mathbf{i}}_{u,2},\ldots,\hat{\mathbf{i}}_{u,m}\}$ and target item embeddings $\mathbf{I}_{u}=\{\mathbf{i}_{u,1},\mathbf{i}_{u,2},\ldots,\mathbf{i}_{u,n}\}$ for user $u$, along with random negative embeddings $\mathbf{N}$, we calculate the proportion of target items for which at least one predicted embedding retrieves it within the top-$k$ results:
$\begin{split}\mathrm{ur}_{k}(\mathbf{\hat{I}},\mathbf{I},\mathbf{N})\coloneq\frac{1}{|\mathbf{\mathbf{\hat{I}}}|}\times\sum_{i=1}^{|\mathbf{\hat{I}}|}\mathbf{1}\left\{\min_{j,z\in[\Delta]}\left|\{\mathbf{n}\in\mathbf{N}\mid\mathbf{\hat{I}}_{ij}^{\top}\mathbf{n}\geq\mathbf{\hat{I}}_{ij}^{\top}\mathbf{I}_{iz})\}\right|\leq k\right\}\end{split}$
where $\text{rank}(\mathbf{i},\hat{\mathbf{i}},\mathbf{N})$ denotes the rank of target item $\mathbf{i}$ when items in $\mathbf{N}\cup{\mathbf{i}}$ are sorted by cosine similarity to prediction $\hat{\mathbf{i}}$. This metric captures whether any of the generated embeddings successfully retrieves each target item, reflecting the practical scenario where multiple predictions can contribute to successful recommendation.
We perform autoregressive generation without teacher forcing, ensuring the transformer is not conditioned on any target items during prediction, making the evaluation more realistic for deployment scenarios.
Beyond retrieval performance, we examine recommendation diversity by computing the proportion of unique items retrieved across all users in the evaluation set after selecting top-$k$ predictions. This quantifies whether learned representations exhibit bias toward popular items versus providing diverse recommendations across the user base.
5.2. Offline Results
In this section, we compare PinRec against existing sequence-based user modeling approaches, including Pinnerformer (Pinnerformer22), SASRec (kang2018self), and the leading generative recommendation method TIGER (Tiger23). We also evaluate our transformer architecture against the proposed HTSU (HSTU24) module. Furthermore, we quantify the impacts of our design choices on both recommendation performance and diversity.
5.2.1. Comparison with Prior Method
We evaluate three different variants of PinRec.
  • PinRec-UC: We do not condition on anything (i.e., $\Phi=\emptyset$) and do not employ multi-token prediction.
  • PinRec-OC: We condition on the target action (“outcome conditioned”) and do not employ multi-token prediction.
  • PinRec-{MT, OC}: We condition on the target action and employ multi-token prediction.
To ensure that the comparison between PinRec-{MT, OC} and other methods is fair, we adjust the number of autoregressive generation steps for multi-token generation to ensure that the total number of generated embeddings is held constant. As in training (where we cannot modify the evaluation dataset post-hoc), we condition on the target item’s action for outcome-conditioned variants. For simplicity, we reproduce results for the outcome-conditioned variant only when comparing against HTSU (additional details in Section A.1.1). The impact of negative sampling and model scaling effects are analyzed in Section A.1.2 and Section A.1.3, respectively.
Our comparison in Table 1 shows that PinRec outperforms other baselines across the board. Importantly, outcome-conditioning demonstrates notable improvements over PinRec-UC (unconditioned) in unordered recall @ 10 (e.g., +2-3% on Homefeed and Related Pins). Above all, PinRec-{MT, OC} shows improvements over every method, with over +10% lift on Homefeed, over +20% on Related Pins, and +30% on search over PinRec-UC.
Table 1. Comparison of baselines and PinRec variants (unordered recall @ 10) across major surfaces at Pinterest.
Model Homefeed Related Pins Search
SASRec (kang2018self) 0.382 0.426 0.142
PinnerFormer (Pinnerformer22) 0.461 0.412 0.257
TIGER (Tiger23) 0.208 0.230 0.090
HTSU-OC (HSTU24) 0.596 0.539 0.179
PinRec-UC 0.608 0.521 0.350
PinRec-OC 0.625 0.537 0.352
PinRec-{MT, OC} 0.676 0.631 0.450
5.2.2. Outcome-Conditioned Generation
We assess whether conditioning on a desired action (e.g., “repin”) leads to outputs that elicit that action. If our recommendations successfully control outcomes, we expect higher offline recall for the conditioned action and lower recall for others. As a baseline, we compare against the unconditioned PinRec variant.
For simplicity, we allocate the entire budget to one action and perform offline evaluation conditioning across each action. We show the result of this analysis in Figure 5. Whenever we condition on an action outcome, we show notable gains over an unconditioned baseline for that action. Specifically, matching outcome conditioning yields 1.9% recall lift for true repins, 1.4% on true grid clicks, 4.8% on long grid clicks, and 6.2% on outbound clicks.
On the other hand, we observe decreases in recall compared to an unconditioned baseline if the true action does not correspond to the desired action, indicating that our conditioning achieves the desired result of producing outcome-specific representations. Especially for outbound clicks, we observe that conditioning on any other actions leads to massive drops in unordered recall.
Figure 5. Percentage lift in unordered recall for PinRec-OC over PinRec-UC when conditioning on the desired action, stratified by the actual action taken by the user.
5.2.3. Utility of Multi-Token Generation
We evaluate the benefits of multi-token generation as a function of the number of generations produced per timestep. Specifically, we analyze the tradeoff between generating more representations in fewer steps versus fewer representations over more steps, while keeping the total number of output embeddings constant. Since the strongest offline performance aside from the multi-token variant is by PinRec-OC, we consider it as a baseline. For simplicity, we average the unordered recall @ 10 across the three surfaces.
We show the tradeoff between performance, diversity, and efficiency with the number of embeddings we generate in each generation step using multi-token generation in Figure 6. While next-token generation (i.e., 1 embedding per step) is the most inefficient and yields the poorest performance (where we perform similarly to PinRec-OC), we observe significant benefits across the board in increasing the number of embeddings per step. At 16 generations per step, we demonstrate a $\sim$10x reduction in latency over PinRec-OC, with improved unordered recall (+16.0%) and diversity (+21.3%).
Figure 6. Tradeoff between unordered recall @ 10 (top), proportion unique (bottom), and generation latency (right) with the number of multi-token generations per step.
6. Online A/B Experiments
We perform online A/B experiments to evaluate PinRec as a candidate generator, measuring its impact on user engagement and confirming its online controllability. The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other candidate generators, to produce the final results shown to users. We tested on the three major surfaces: Homefeed, Search and Related Pins.
6.1. Homefeed
6.1.1. Experiment Configuration
In the Homefeed online experiment, we run five groups exposed to global traffic for PinRec-OC, with varying action budgets $\mathcal{B}$. The budgets are fixed for outbound click (0.1), long grid clicks (0.05), and other unspecified actions (0.05). The remaining budget is split between the repin and grid click engagements, in steps of 0.2 ranging from 0 to 0.8. We similarly configure five groups for PinRec-{OC, MT}, with the same budgets. We include control groups without PinRec.
6.1.2. Metrics
We use engagement actions such as repins, grid clicks, and time spent. Additionally, we count “fulfilled sessions,” where the user takes at least one action in the session, and “unfulfilled sessions,” where the user does not take any action.
6.1.3. Results
In Figure 7, we visualize a Pareto front of the repin rate (relative to impressions) and grid click rate for the PinRec candidate generator by itself, with absolute numbers hidden due to non-public information. By reducing the grid click budget and correspondingly increasing the repin budget, we see the expected result as the grid click rate drops and repin rate increases. Similarly, we observe a significant correlation between the budgets and the overall A/B experiment metric lifts for Homefeed grid clicks and repins in Figure 8 in Appendix.
(a) PinRec-OC
(b) PinRec-{MT, OC}
Figure 7. Visualization of Pareto fronts for online rate of grid click versus repin for items recommended only by PinRec variants, with error bars denoted in grey.
For each PinRec variant, we selected the variant with the best fulfilled sessions and measured the percentage lift in metrics across Homefeed and the entire site (Table 2). Relative to the control group, the unconditioned variant of PinRec, PinRec-UC, demonstrates improvements in unfulfilled sessions and in Homefeed and site-wide grid clicks. In addition to those lifts, the outcome-conditioned variant, PinRec-OC, shows significant lift in site-wide fulfilled sessions. The multi-token variant, PinRec-{MT, OC}, significantly increased successful site sessions, time spent on site, and site-wide grid clicks.
Table 2. A/B experiments for PinRec variants as candidate generators in Homefeed.
Type Metric UC OC {MT, OC}
Overall Fulfilled Sessions +0.02% +0.21% +0.28%
Metrics Time Spent -0.02% +0.16% +0.55%
Unfulfilled Sessions -0.28% -0.89% -0.78%
Action Site-wide Grid Clicks +0.58% +1.76% +1.73%
Metrics Homefeed Grid Clicks +1.87% +4.01% +3.33%
6.2. Search
On the Search surface, we evaluate the unconditioned variant of PinRec. As shown in Table 3, PinRec-UC improves search metrics significantly relative to a control group, with increased session-wide search fulfillment (sessions resulting in positive interactions like repins or long clicks). When deployed for promoted product retrieval, the same embeddings achieve a 1.83% reduction in CPA and 1.87% increase in shopping conversion volume in the US, indicating improved relevance and user engagement with sponsored content.
Table 3. A/B experiments for PinRec as candidate generator in Search retrieval.
Metric Type Metric PinRec-UC
Overall Metrics Search Fulfillment Rate +2.27%
Action Metrics Search Repins +4.21%
Search Grid Clicks +3.00%
6.3. Related Pins
On the Related Pins surface, we evaluate the outcome-conditioned variant of PinRec. As shown in Table 4, PinRec-OC improves Related Pins metrics significantly relative to a control group. We observe session-wide increase in fulfilled sessions and increase in sitewide engagement.
Table 4. A/B experiments for PinRec variants as candidate generators in Related Pins.
Type Metric PinRec-OC
Overall Fulfilled Sessions +0.26%
Metrics Time Spent +0.30%
Unfulfilled Sessions -1.07%
Action Site-wide Grid Clicks +0.50%
Metrics Sitewide Repins +0.63%
7. Conclusion
We introduced PinRec, a generative retrieval system developed for Pinterest’s Homefeed, Search and Related Pins feed. To the best of our knowledge, this paper represents the first comprehensive study on deploying generative retrieval in a web-scale application. The incorporation of two key innovations — outcome-conditioned generation and temporal multi-token inference — enabled PinRec to enhance several business metrics without substantial cost increases.
Several promising future directions can be pursued. First, applying generative approaches to ranking, as explored by Zhai et al. (HSTU24), could be beneficial. Second, optimizing input sequences to the model, akin to prompt engineering in large language models, is another avenue worth exploring. Third, developing language model alignment techniques similar to Ouyang et al. (PPO22) could enhance the use of engagement signals from users. Overall, this paper highlights a promising avenue for utilizing generative retrieval in large-scale recommender systems.
Appendix A Appendix
A.1. Additional Analysis
In this section, we perform additional experiments to quantitatively and qualitatively show that the design decisions made for PinRec are justified and that it performs across the state-of-the-art architectures. Additionally, we provide more online analyses of the performance across different user cohorts and tail items.
A.1.1. HSTU Architecture
Though PinRec relies primarily on the well-known transformer architecture, we note that our technique is completely architecture-agnostic. Specifically, outcome conditioning and multi-token prediction can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements. To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (HSTU24), as opposed to the multi-head attention-based causal transformer. Note that we do not leverage the ranking task or any sequence sampling explored in (HSTU24) as it is not relevant for our specific use case. For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
The results are shown in Table 5, which clearly shows that HSTU is roughly 3% worse on HomeFeed (effectively negating the improvement with respect to the unconditioned transformer-based PinRec) and competitive on the related pins surface. However, there is a substantial drop in performance on the search surface, which is a critical component of the user experience at Pinterest. We believe that this may be related to specific architectural elements of HSTU.
Table 5. Comparison of HSTU and transformer architectures for PinRec, where PinRec-OC leverages the transformer as presented in the main text.
Model Homefeed Related Pins Search
PinRec-OC 0.625 0.537 0.352
HSTU (OC) 0.596 0.539 0.179
Though we observe a sharp drop in the search surface, HSTU performs roughly on the same order of magnitude as the traditional transformer architecture, demonstrating that our ideas perform across different architectures. This validates that our ideas are robust to architectures across the board.
A.1.2. Ablating Negative Sampling
Following from Pinnerformer22, we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function. Though we ultimately end up with similar conclusions as Pinnerformer22, we show these findings for maximum clarity regarding our design choices.
Random Negatives
Though in-batch negatives are a critical component of our negative corpus, we construct a random negative corpus consisting of a random set of possible items. Given that our retrieval problem is ultimately across a random selection of pins, we wish for PinRec to show retrieval capability across this corpus. We perform several ablations with random negatives, with in-batch negatives constant at a non-zero value. For these analyses, we focus on Homefeed (the most critical surface) and present results aggregated across all surfaces as well.
We show results in Table 6, where there are moderate gains in unordered recall if there is a non-zero number of in-batch negatives present. Specifically, on the most critical surface of Homefeed, we see notable improvement over the baseline in terms of unordered recall.
Table 6. Percentage lift in unordered recall using 16K random negatives, relative to using only 8K random negatives.
Surface % Lift in Recall @ 10
Homefeed +2.5%
All +1.3%
To summarize, while random negatives are certainly an important piece of the puzzle, they cannot substitute in-batch negatives. When used in conjunction, scaling the number of negatives yields notable improvement, especially on Homefeed.
In-Batch Negatives
As shown in the previous subsection, disabling in-batch negatives entirely leads to significantly poorer metrics, showing their importance. Can we improve performance offline by scaling the batch size and consequently, the number of in-batch negatives? To examine this, we fix the number of random negatives and ablate the number of in-batch negatives through the batch size.
Table 7. Percentage lift in unordered recall @ 10 through scaling the number of in-batch negatives (relative to baseline of 1x) with random negatives fixed.
# In-Batch Negs. % Lift (HF) % Lift (all)
1.5x +4.2% +3.3%
2x +8.4% +5.4%
12x +18.3% +12.1%
We show results in Table 7, which demonstrates a clear trend that shows that a greater number of in-batch negatives and batch size leads to better performance, especially for the Homefeed surface. These are similar in directional improvements to random negatives, but we observe greater improvements with in-batch negatives, which are likely more relevant as negative examples.
A.1.3. Ablating Parameter Scaling
One of the recent explorations within recommendation systems has been scaling to improve performance, in an attempt to replicate the success of these efforts for language modeling. To explore this hypothesis, we scale the number of parameters within PinRec across two fronts: the “vocabulary” and the model itself. While model scaling is clear and involves scaling up the embedding size, the number of transformer layers, and so on, scaling the vocabulary for an embedding-based generative retrieval technique involves improving the input representations. Specifically, as mentioned in the main text, we implement “ID” embeddings for pins, that effectively serve as additional representation for the “pin vocabulary”.
Model Parameters
As presented in the main text, PinRec (all variants) have roughly 100 million transformer parameters, corresponding to the base form or size of GPT-2. To ablate the effect of the size of the transformer, we attempt to follow the different scales employed by GPT-2 (base, medium, large, XL) as they have shown success in terms of the factors by which the embedding size, number of layers, and number of heads have scaled. Note that due to computational constraints, we reduce the scaling in terms of the number of layers, which increases GPU memory footprint significantly. Note that we present lift in offline unordered recall @ 10 for the outcome conditioned variant of PinRec below, but we observe similar trends for the unconditioned model as well.
Table 8. Percentage lift in unordered recall through expanding the transformer at various GPT-2 scales relative to base ( 100 million parameters).
Model % Lift in Recall @ 10
Medium -0.8%
Large -0.9%
XL +2.2%
Though we observe some gains in Table 8 from scaling the number of transformer parameters to around 1 billion (XL-sized), these become intractable to serve in real time, requiring offline batch inference or significantly higher cost and more optimization. For the remaining sizes, we observe insignificant lifts (note that less than 1% in lift magnitude can be often explained through noise). Hence, we opt to not use larger sized transformers.
Vocabulary Parameters.
As presented in the main text, the ID embeddings consist of roughly 10 billion parameters. These are a core component of the model since they provide additional representational power for the most important entity type at Pinterest: Pins. To ablate the importance of these ID embeddings, we examine the performance in offline recall across all surfaces with and without ID embeddings. For this analysis, we focus on Homefeed (the most critical surface) and present results aggregated across all surfaces as well in Table 9.
Table 9. Percentage lift through adding ID embeddings (10 billion parameters) over no ID embedding-based vocabulary.
Surface % Lift in Recall @ 10
Homefeed +14.0%
All +14.4%
Evidently, the improvement in recall is substantial and indicates that the scaling of vocabulary to learn an embedding for each unique ID (roughly, as it is hash-based) rather than only content-based features is beneficial.
A.1.4. Tail User and Item Analysis
One of the potential drawbacks of outcome conditioning would be for “tail” or “fresh” users, i.e., those which have lesser activity or less propensity for high-effort engagement. Critically, our method intends to optimize for outcomes across all cohorts of users, including such users. In the same vein, we wish to optimize for “tail” items, which may receive a lesser number of impressions, in an effort to distribute content to improve corpus coverage.
User Cohort Grid Clicks Outbound CTs
UC OC UC OC
New +2.60% +2.62% +3.98% +5.78%
Resurrected +1.56% +1.86% +1.98% +1.72%
(a) Online lift (%) for grid clicks and outbound clickthroughers for new and resurrected users.
Metric UC OC
Corpus Coverage (unique pins) +6.2% +17.3%
1-Imp Pin Rec. (tail items) +5.8% +7.7%
(b) Relative lift (vs. 2-tower model) for tail item and corpus coverage metrics.
Table 10. Online experimental results for different user cohorts and tail item metrics, across PinRec-UC and PinRec-OC.
In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 10(a). Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users. In fact, across new users, OC performs better on outbound CTs.
To examine performance for tail items, we compare to a production 2-tower model in an online setting in Table 10. On corpus coverage, adding PinRec-UC/OC surfaces 6.2%/17.3% more unique pins relative to the 2-tower. For long-tail items with 1-impression, PinRec-UC/OC achieve 5.8%/7.7% absolute lifts respectively, relative to the aforementioned 2-tower model. Evidently, across the board, PinRec surfaces more of the corpus and surfaces more tail items compared to control, with outcome conditioning showing even more significant lifts over control in both regards.
We believe the evidence suggests that our generative retrieval solution is equal to or better across inactive/new users and tail items relative to existing control, and perhaps more importantly, it shows that outcome conditioning can help in boosting performance in these criteria, which is one of its goals.
(a) PinRec-OC
(b) PinRec-{MT, OC}
Figure 8. Visualization of Pareto fronts for online lift % in Homefeed grid clicks versus repins for PinRec variants.
A.1.5. Serving Efficiency
One of the critical aspects of our proposed generative retrieval technique is an effort to maintain serving efficiency. With our custom infrastructure built from off-the-shelf serving technologies such as NVIDIA Triton Inference Server and optimization techniques like torch.compile, we achieve a system that can be served in real-time. In this section, we delve deeper into the memory and time requirements, comparing our method with prior methods deployed in production.
On the memory front, all variants of PinRec (unconditioned, outcome-conditioned with or without multi-token generation) use about 1.6GiB of GPU memory including KVCache and CUDA graphs, without any significant memory usage increase for multi-token prediction or outcome conditioning. Excluding the KVCache and CUDA graphs, we use roughly the same amount of memory (2̃50 MiB) relative to other techniques (i.e., two-tower models).
In terms of latency, at 80 QPS, the online and offline latency for PinRec-OC/OC, MT are p50 40ms/p90 65ms (6 steps). On the other hand, across only 3 steps for PinRec-UC, we observe a latency of p50 23ms and p90 46ms (note that this produces significantly less embeddings and less compressed embeddings than either of the above techniques, though it is lower latency). Though the latency is roughly 3-4x higher (p50) than a traditional two-tower solution, we note that parallelizing it with other RPCs results in ¡1% added end-to-end increase in latency.
A.2. Visualization of Multi-Token Impact
(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
(b) Results for PinRec-OC candidate generator for user history in subfigure above. Notably only includes Pins related to nails, the user’s most recent interaction.
(c) Results for PinRec-{MT, OC} candidate generator for user history in subfigure above. Includes a mix of Pins related to both tattoos and nails.
Figure 9. Comparison of results for a manually-specified user history.
(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
(b) Results for PinRec-OC candidate generator for given user history. Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
(c) Results for PinRec-{MT, OC} candidate generator for given user history. Includes a mix of Pins related to nurseries and food.
Figure 10. Comparison of results for another manually-specified user history.
In this section, we present additional qualitative examples through visualizations of retrieved examples for sample user histories, comparing multi-token and non-multi-token variants. Through these visualizations, we are able to demonstrate a reasonable and consistent qualitative difference in the types of pins retrieved, justifying the improved session-wide and sitewide online metrics for PinRec-{MT, OC}.
Given a user history, we display two examples of retrieved pins in Figure 9 and Figure 10 for PinRec-{MT, OC} and PinRec-{OC}. Specifically, note that PinRec-{MT, OC} captures less-recent interactions for users compared to PinRec-OC. This is expected as the multi-token model predicts over a longer future time window and is less biased toward recency. PinRec-{MT, OC} can therefore produce more diverse results capturing more medium-term interests. On the other hand, since the PinRec-{OC} is naturally biased towards the short term, it fails to capture potential longer term trends that may induce engagement.
Note that these trends reflect the same takeaways from offline and online evaluation. Offline, we observe an improvement in diversity as the number of multi-token predictions is increased, which is consistently above the non-multi-token variant of PinRec. Online, we see that the “feed” presented by PinRec-{MT, OC} is of higher quality due to its diversity and visual appeal/aesthetic. These are critical qualities for a platform like Pinterest.